OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 120r3 square inches. Each tile has an area of the square root of the quantity 16 times r to the ninth power.. Use the given information to write and simplify an expression for the total number of tiles used. I need help

OpenStudy (anonymous):

Number of tiles = area of floor ÷ area of each tile.

OpenStudy (anonymous):

So the numbers that are there I plug them in?

OpenStudy (anonymous):

Yep. Is that 120r3 mean 120r^3?

OpenStudy (anonymous):

Yes thats what it means

OpenStudy (anonymous):

You can simplify this?\[\large n=120r^3 \div \sqrt{16r^9}\]

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

Here's a big hint: \[\large r^9=(r^3)^2\]

OpenStudy (anonymous):

Ohh ok I think I do